OpenStudy (anonymous):

Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y? A. 18 B. 12 C. 6 D. 3 Need help !

OpenStudy (anonymous):

6 = y

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I need it to be explained. how you got the answer. i dont really get the answer

OpenStudy (mathmale):

Interest = Principal * rate * time. Please, Bree, look up the formula for simple interest (given above). Have you a texbook available? Could you do an Internet search (e. g., for "formula for simple interest"? The more you're involved in this problem solving, the more you'll understand about the process.

OpenStudy (anonymous):

I=prt plug in your percentage and amount invested and your interest 612 = 0.085 * 1200 * t 612/102 = 102t /102 6 = t

OpenStudy (mathmale):

@crabby_patty : Nice work. But please provide just enough info so that Bree can find the answer herself; don't give away answers. Thank you for understanding.